A firm has the following total cost function 𝑻π‘ͺ = πŸπŸŽπŸŽπ’’ βˆ’ πŸ“π’’2 + 𝟎. πŸ“π’’3.


a. Find the average cost function.


1
Expert's answer
2022-04-14T15:44:18-0400

Firstly find the MC. It will be MC = "− 10 \cdot𝒒 + 1. 𝟓 \cdot𝒒^2."

Than we make MC = 0, so "− 10 \cdot q + 1. 5 \cdot q^2 = 0
."

q = 6,6. So it was found equilibrium q.

Then use formula AC = TC/q.

TC = 585

Q = 6,6

AC = 88,78